# Argument Rarity-based Originality Assessment for AI-Assisted Writing

### v1

@@ Abstract
As Large Language Models (LLMs) have become capable of effortlessly generating high-quality text, traditional quality-focused writing assessment is losing its significance. If the essential goal of education is to foster critical thinking and original perspectives, assessment must also shift its paradigm from quality to originality. This study proposes Argument Rarity-based Originality Assessment (AROA), a framework for automatically evaluating argumentative originality in student essays. AROA defines originality as rarity within a reference corpus and evaluates it through four complementary components: structural rarity, claim rarity, evidence rarity, and cognitive depth. The framework quantifies the rarity of each component using density estimation and integrates them with a quality adjustment mechanism, thereby treating quality and originality as independent evaluation axes. Experiments using human essays and AI-generated essays revealed a strong negative correlation between quality and claim rarity, demonstrating a quality-originality trade-off where higher-quality texts tend to rely on typical claim patterns. Furthermore, while AI essays achieved comparable levels of structural complexity to human essays, their claim rarity was substantially lower than that of humans, indicating that LLMs can reproduce the form of argumentation but have limitations in the originality of content.
@@
journal: Computers Education: Artificial Intelligence
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) is fundamentally transforming the nature of human intellectual activities [8, 35]. LLMs, exemplified by ChatGPT, have acquired the ability to instantly generate logically coherent, fluent, and persuasive text on any given topic. The capability to write well, once the exclusive domain of humans, is no longer unique to human beings. While this technological innovation is affecting various domains of society, its impact is particularly profound in educational settings. Determining whether an essay or report submitted by a student is truly the product of original thinking or merely a reproduction of typical patterns has become more challenging than ever before.
@@
Looking back at the history of writing assessment, evaluation criteria have consistently emphasized quality. Logical coherence, grammatical accuracy, clarity of organization, and persuasiveness of argumentation have been widely recognized as universal characteristics of excellent writing. Research on Automated Essay Scoring (AES) has primarily focused on predicting these quality indicators [18]. However, now that LLMs have reached a level where they surpass humans in quality, assessment based solely on such criteria is entering a stage where its fundamental significance must be reconsidered. If AI can generate essays of near-perfect quality, high quality alone is no longer a sufficient indicator reflecting human thinking ability or learning outcomes.
@@
What should be questioned here is the essential purpose of having students write in education. The true purpose of assigning essays to students is not the production of well-formed text itself. Rather, it lies in having students experience the process of critical thinking: thinking deeply about a given topic, forming original perspectives, exploring diverse evidence, and anticipating and responding to counterarguments [13]. In other words, what education should originally evaluate is not the quality of writing but the originality of the thinking behind it. From this perspective, the paradigm of assessment must shift from quality to originality.
What then does argumentative originality specifically refer to? This study conceptualizes originality not simply as being different from others in a superficial sense, but as multifaceted characteristics of argumentation. First, there is the uniqueness of the argument structure itself. This dimension concerns whether the structural pattern of how claims are supported by reasons, how evidence is arranged, and how counterarguments are addressed is typical or rare. Second, there is the uniqueness of claim content. This dimension concerns whether the claim is one that many people would make on the same topic or whether it represents an original perspective. Third, there is the uniqueness of evidence selection. This dimension concerns whether the argument relies on commonly cited typical evidence or employs rare evidence that was independently explored. Fourth, there is the depth of reasoning. This dimension concerns whether the argument remains at a simple claim-evidence correspondence or develops multi-layered reasoning that anticipates counterarguments and constructs rebuttals. These dimensions are mutually independent, and being original in one dimension does not guarantee originality in others. For truly comprehensive originality assessment, these multifaceted aspects must be captured individually and integrated.
However, consistent evaluation of such argumentative originality by human raters involves inherent difficulties. In educational settings, instructors must grade dozens to hundreds of essays within limited time at the end of a semester. Under such circumstances, it is difficult to avoid order effects and fatigue in grading, where impressions from earlier essays unconsciously influence subsequent evaluations [19]. Additionally, the evaluator’s own expertise and prior knowledge act as biases [1]. Evaluators tend to give higher scores to claims that align with their expectations and beliefs, while finding it difficult to appropriately recognize the value of unfamiliar perspectives or rare evidence. As an even more fundamental problem, judging how rare an essay’s claims or evidence are within the entire corpus when evaluating individual essays exceeds the limits of human cognitive capacity. It is practically impossible for evaluators to simultaneously remember the content of hundreds of essays and judge relative rarity. Due to these limitations, originality assessment tends to be subjective and inconsistent, making it difficult to provide fair and meaningful feedback to students.
In recent years, new possibilities have emerged through leveraging the advanced language understanding capabilities of LLMs. LLMs possess the ability to deeply understand the semantic content of text, analyze argument structure, and grasp relationships between elements [20]. By applying this capability to assessment, there is potential to complement the limitations of human evaluators. LLMs do not fatigue, are not affected by order effects, and can evaluate all essays according to the same criteria. They can also process large-scale corpora and objectively calculate the relative positioning of each essay [3]. Indeed, such AI-assisted assessment is already being implemented in practice. For example, at Ritsumeikan University, a large private university in Japan, on-demand data science courses accommodate up to 1,000 students per class. To address this scale, the authors have independently developed and deployed an AI-powered assessment system called RaaS (Ritsumeikan AI-powered Assessment Solution), which is already operational in actual courses[25]. While final evaluation responsibility remains with human instructors, this represents an emerging case of large-scale course operation premised on AI-assisted assessment. Despite such practical implementations, most existing research on AI-assisted assessment has focused on quality evaluation and grammatical error detection, and no framework has been established for systematically evaluating originality. The framework for quality assessment cannot be directly applied to originality assessment. Originality is an essentially different concept from quality and requires its own theoretical foundation and evaluation methodology.
This study proposes Argument Rarity-based Originality Assessment (AROA), a framework for automatically evaluating argumentative originality in argumentative essays. The core idea of AROA is to operationally define originality as rarity within a reference corpus. For a claim in an essay to be original means that the claim is rare among other essays addressing the same topic. This definition makes it possible to measure originality, an abstract concept, as a quantitative indicator based on density estimation.
@@
AROA adopts a multifaceted approach that conceptualizes originality as four complementary components. The first component, structural rarity, quantifies how rare the pattern of argument structure consisting of claims, reasons, evidence, counterarguments, and rebuttals is within the reference corpus, using kernel density estimation. The second component, claim rarity, measures how far the central claims of an essay are from other essays in semantic space, using neighborhood density of sentence embeddings. The third component, evidence rarity, evaluates the uniqueness of evidence used to support claims using the same methodology. The fourth component, cognitive depth, quantifies the dialogic structure of counterargument-rebuttal exchanges and the hierarchical depth of reasoning as complexity of cognitive operations. These four components have been confirmed through correlation analysis to have low mutual correlations, each capturing different aspects of originality. AROA integrates these components after adjusting them with quality scores, thereby treating quality and originality as independent axes while calculating a comprehensive originality score that considers the balance between both.
The contributions of this study are summarized in the following three points.
  1. (i)
@@
    AROA is a novel framework designed to automatically evaluate argumentative originality in argumentative essays through four complementary components—structural rarity, claim rarity, evidence rarity, and cognitive depth—integrating density-based rarity quantification with quality adjustment.
  2. (ii)
@@
    Empirical demonstration of a strong negative correlation between text quality and semantic rarity through large-scale experiments using human essays and AI-generated essays, revealing a quality-originality trade-off where logically coherent, high-quality texts tend to rely on typical claim patterns and common evidence.
  3. (iii)
@@
    Comparative analysis of component profiles between human essays and AI essays, revealing that contemporary LLMs achieve comparable levels to humans in structural complexity of argumentation while exhibiting notable limitations in semantic originality of claims and evidence, thereby presenting new guidelines for originality assessment in the AI era.
@@
The remainder of this paper is organized as follows. Section 2 reviews related work and clarifies the positioning of this study. Section 3 describes the details of the proposed framework AROA. Section 4 presents the experiments and results, and Section 5 discusses the implications and limitations of the results. Finally, Section 6 presents the conclusions.
@@ 2.1 Automated Essay Scoring
AES is a technology that automatically assigns scores to essays using computers. From the late 1990s to the 2000s, practical systems such as e-rater [7], IntelliMetric, and Intelligent Essay Assessor were developed and deployed in large-scale standardized tests such as GMAT and GRE. Early AES systems extracted handcrafted features such as lexical diversity, syntactic complexity, and discourse structure, and predicted scores using regression models. However, Powers et al. [24] conducted experiments in which professional writers intentionally created essays to deceive e-rater, demonstrating that texts that were content-poor but satisfied formal features tended to receive high scores. This study made it clear that AES at the time relied heavily on surface features and could not adequately evaluate the validity of logic or depth of content.
Since 2016, with the advancement of deep learning, the technical paradigm of AES has undergone a major transformation. Taghipour and Ng [30] proposed an end-to-end AES using LSTM and demonstrated the effectiveness of neural approaches that eliminate the need for handcrafted feature engineering. Hussein et al. [15] provided a comprehensive review that systematically classified AES research into handcrafted feature-based and automatic feature learning-based approaches, identifying three major challenges of AES: human-likeness, deception resistance, and creativity assessment.
In recent years, the application of large language models (LLMs) such as GPT to AES has rapidly progressed. Naismith et al. [20] proposed a method to evaluate discourse coherence in essays using GPT-4 and demonstrated the possibility of explainable AES that can also generate evaluation rationales. Xiao et al. [36] demonstrated that novice evaluators who referred to LLM feedback achieved accuracy close to that of experts, presenting a perspective that redefines AES not as complete automation but as human capability augmentation. However, Bevilacqua et al. [4] reported that LLM-based AES tends to rate GPT-generated texts higher than human texts, clarifying a fundamental bias problem of AES in the generative AI era.
Surveying these prior studies reveals that AES research has consistently focused on predicting text quality—logical coherence, grammatical accuracy, and clarity of organization. From early handcrafted feature-based systems to the latest LLM-based approaches, the evaluation target has been whether text is well-written, not whether argumentation is original. As Dikli [10] pointed out, creativity and critical thinking assessment have been recognized as challenges for AES, yet no framework has been proposed to systematically address this. AROA, proposed in this study, addresses this research gap by introducing argumentative originality as an evaluation axis independent of quality, presenting a new direction for AES research.
@@ 2.2 Argumentation Mining
Argumentation Mining is a research field that automatically identifies and extracts argument structures from natural language text. Lawrence and Reed [16] provided a comprehensive survey of this field, systematically organizing topics from the identification of claims and premises to the extraction of support and rebuttal relations, and the mapping to the Toulmin model and argumentation schemes. The argumentation model proposed by Toulmin [31] is widely adopted as the theoretical foundation of argumentation mining. This model structures argumentation through six elements—Claim, Grounds (Data), Warrant, Backing, Qualifier, and Rebuttal—and has become the standard framework for argumentative essay analysis in the educational domain.
Argumentation mining research targeting student essays was formally initiated by Stab and Gurevych [29]. They proposed an annotation scheme for claims, premises, and support/attack relations in persuasive essays, achieving high inter-annotator agreement. This study became the starting point for subsequent research on argument structure analysis in student essays. Wachsmuth et al. [33] proposed a method to evaluate argument quality (organization, strength) in essays using extracted argument structures, extending argumentation mining from structure extraction to quality assessment.
Educational applications of argumentation mining have also progressed. Nguyen and Litman [21] integrated argument features such as the number of claims and support relations into AES and demonstrated improved accuracy in persuasive essay scoring. Wambsganss et al. [34] developed an educational support system that visualizes argument structures and provides feedback to students, showing that pointing out missing claims or insufficient evidence promotes learners’ argumentative self-correction.
In recent years, argumentation mining research using LLMs has rapidly progressed. Naismith et al. [20] proposed a method to evaluate discourse coherence in essays using GPT-4 and demonstrated that it can better reflect argument structure quality than conventional surface cohesion indicators. Li et al. [17] proposed a multi-task benchmark integrating claim detection, evidence extraction, and persuasiveness evaluation, revealing that even LLMs with high performance on single tasks show significant performance degradation when performing argument understanding in a chained manner.
These prior studies have made important advances in argument structure extraction and quality assessment. However, the quality evaluated in existing research concerns aspects such as structural completeness, organization, and persuasiveness, and does not include the originality of argument content, that is, how rare claims and evidence are compared to other essays. AROA, proposed in this study, utilizes argumentation mining techniques to extract claims, evidence, counterclaims, and rebuttals, and then quantifies their semantic rarity through density estimation. This presents a new direction of evaluating originality rather than quality of argument structures.
@@ 2.3 Originality and Creativity Assessment
Originality and creativity assessment has a long tradition of research in psychology and education. Runco and Yoruk [27] defined creativity as the combination of novelty and usefulness, and demonstrated that divergent thinking is the representative paradigm for measuring this. Among divergent thinking tests, the Alternative Uses Test (AUT) is the most widely used, asking respondents to generate as many alternative uses as possible for a given object, and evaluating responses in terms of fluency, flexibility, originality, and elaboration.
Traditionally, originality assessment in AUT was conducted by human raters subjectively judging the rarity of responses, but in recent years, automatic evaluation based on semantic distance has become mainstream. Beaty et al. [2] proposed a method to automatically evaluate originality by calculating the semantic distance between stimulus words and response words using sentence embeddings. This method showed moderate to high correlation with human evaluation and was established as the standard approach for automatic creativity assessment. Patterson et al. [23] verified the multilingual validity of semantic distance-based originality assessment using large-scale data from 12 languages and over 6,500 participants, demonstrating that Transformer-based embeddings enable stable evaluation across languages.
In recent years, creativity assessment using LLMs has rapidly developed. DiStefano et al. [9] fine-tuned LLMs for the higher-order creative language task of metaphor generation and achieved agreement with human evaluation that significantly exceeded semantic distance methods. Valueva et al. [32] comprehensively reviewed the evolution of automatic scoring methods for divergent thinking tests, pointing out that while LLMs are approaching human evaluation, the opacity of judgment rationales is a new challenge. On the other hand, Runco et al. [26] analyzed divergent thinking responses generated by GPT-3.5 and GPT-4, showing that while AI can generate numerous ideas, it also includes many irrelevant and redundant expressions. This result reconfirms the importance of the definition that creativity equals novelty plus usefulness.
These prior studies have succeeded in evaluating the novelty of short responses such as those in divergent thinking tests. However, no research exists that evaluates argumentative originality in long texts such as argumentative essays, that is, how rare argument elements such as claims, evidence, and counterarguments are. Existing creativity assessment research measures semantic distance of words or short sentences, and no framework has been proposed to capture the originality of entire argument structures in a multifaceted manner. AROA, proposed in this study, integrates the concept of rarity-based evaluation with argumentation mining, assessing argumentative originality through four components: claim rarity, evidence rarity, structural rarity, and cognitive depth. This presents a new direction that extends the concept of rarity as originality, established in divergent thinking research, to complex texts such as argumentative essays.

### v3

@@ Abstract
This study proposes Argument Rarity-based Originality Assessment (AROA), a framework for automatically evaluating argumentative originality in student essays. AROA defines originality as rarity within a reference corpus and evaluates it through four complementary components: structural rarity, claim rarity, evidence rarity, and cognitive depth, quantified via density estimation and integrated with quality adjustment. Experiments using 1,375 human essays and 1,000 AI-generated essays on two argumentative topics revealed three key findings. First, a strong negative correlation (r=0.67r=-0.67) between text quality and claim rarity demonstrates a quality-originality trade-off. Second, while AI essays achieved near-perfect quality scores ($Q=0.998$), their claim rarity was approximately one-fifth of human levels (AI: 0.037, human: 0.170), indicating that LLMs can reproduce argumentative structure but not semantic originality. Third, the four components showed low mutual correlations (r=0.06r=0.060.130.13 between structural and semantic dimensions), confirming that they capture genuinely independent aspects of originality. These results suggest that writing assessment in the AI era must shift from quality to originality.
@@
journal: Information Processing & Management
@@ 1.1 Background and Motivation
The rapid advancement of Large Language Models (LLMs) is fundamentally transforming human intellectual activities [Chang et al., 2024, Wang et al., 2024]. LLMs such as ChatGPT can instantly generate logically coherent, fluent, and persuasive text on any topic. Writing well is no longer an exclusively human capability, which has a particularly profound impact on education, where determining whether a student’s essay reflects original thinking or merely reproduces typical patterns has become increasingly difficult.
@@
Historically, writing assessment has emphasized quality, logical coherence, grammatical accuracy, organizational clarity, and persuasive argumentation. Automated Essay Scoring (AES) research has primarily focused on predicting these quality indicators [Li and Ng, 2024]. However, now that LLMs can surpass humans in text quality, such criteria alone are no longer sufficient indicators of human thinking ability or learning outcomes.
@@
The essential purpose of having students write is not the production of well-formed text itself but the cultivation of critical thinking: forming original perspectives, exploring diverse evidence, and anticipating counterarguments [Hickman, 2023]. From this perspective, the paradigm of assessment must shift from quality to originality.
This study conceptualizes argumentative originality as four mutually independent dimensions: i) the uniqueness of argument structure: whether the pattern of claims, reasons, evidence, and counterarguments is typical or rare; ii) the uniqueness of claim content: whether the central assertion represents an original perspective; iii) the uniqueness of evidence selection: whether the argument employs rare, independently explored evidence; and iv) the depth of reasoning: whether the argument develops multi-layered reasoning with counterargument-rebuttal exchanges. These dimensions are mutually independent, and comprehensive originality assessment requires capturing each individually.
However, consistent evaluation of such originality by human raters is inherently difficult. Instructors face order effects and fatigue when grading large numbers of essays [Mahshanian and Shahnazari, 2020], their own expertise introduces bias [Alvén, 2019], and judging relative rarity across an entire corpus exceeds human cognitive capacity. LLMs offer a promising alternative: they are immune to fatigue and order effects, can evaluate essays according to uniform criteria, and can process large-scale corpora to objectively calculate relative positioning [Naismith et al., 2023, Beaty and Johnson, 2021]. Indeed, AI-assisted assessment is already being implemented in practice; for example, the RaaS system deployed at a large Japanese university supports courses with up to 1,000 students per class [Ritsumeikan University Institute for Teaching and Learning, 2025]. Despite such implementations, most existing AI-assisted assessment research has focused on quality evaluation; no framework has been established for systematically evaluating originality, which is an essentially different concept from quality and requires its own theoretical foundation.
@@
This study proposes Argument Rarity-based Originality Assessment (AROA), a framework for automatically evaluating argumentative originality. The core idea of AROA is to operationalize originality as rarity within a reference corpus and to measure it quantitatively using density estimation. AROA evaluates originality through four complementary components: structural rarity (atypicality of argument structure via kernel density estimation), claim rarity (semantic distance of central claims via neighborhood density of sentence embeddings), evidence rarity (uniqueness of supporting evidence using the same methodology), and cognitive depth (dialogic structure and hierarchical depth of reasoning). These components have been confirmed through correlation analysis to have low mutual correlations, each capturing different aspects of originality. AROA integrates them after adjusting with quality scores, thereby treating quality and originality as independent evaluation axes.
1.2 Research Objectives
Based on the above background and motivation, this study addresses the following three research questions:
  1. RQ1:
    Can argumentative originality in essays be automatically evaluated in a multifaceted manner through density-based rarity quantification of argument components?
  2. RQ2:
    What is the relationship between text quality and argumentative originality, and does a trade-off exist between them?
  3. RQ3:
    How do the originality profiles of human essays and AI-generated essays differ across structural and semantic dimensions?
To address these research questions, this study proposes the AROA framework and conducts large-scale experiments using 1,375 human essays and 1,000 AI-generated essays. The contributions are summarized as follows:
  1. i)
@@ 3. RQ3:
    AROA is a novel framework that automatically evaluates argumentative originality through four complementary components—structural rarity, claim rarity, evidence rarity, and cognitive depth—integrating density-based rarity quantification with quality adjustment.
  2. ii)
@@
    Empirical demonstration of a strong negative correlation (r=0.67r=-0.67) between text quality and claim rarity through large-scale experiments, revealing a quality-originality trade-off where logically coherent, high-quality texts tend to rely on typical claim patterns and common evidence.
  3. iii)
@@
    Comparative analysis revealing that contemporary LLMs achieve comparable structural complexity to humans in argumentation while exhibiting substantially lower semantic originality of claims (approximately one-fifth of human levels), thereby presenting new guidelines for originality assessment in the AI era.
@@
The remainder of this paper is organized as follows. Section 2 reviews related work and clarifies the positioning of this study. Section 3 describes the proposed framework AROA. Section 4 presents experiments and results. Section 5 discusses implications and limitations. Finally, Section 6 presents the conclusions.
@@ 2.1 Automated Essay Scoring
AES is a technology that automatically assigns scores to essays using computers. From the late 1990s to the 2000s, systems such as e-rater [Burstein et al., 1998], IntelliMetric, and Intelligent Essay Assessor were deployed in large-scale standardized tests. Early systems extracted handcrafted features such as lexical diversity and syntactic complexity, and predicted scores using regression models. However, Powers et al. [2002] demonstrated that content-poor texts satisfying formal features could receive high scores, revealing that early AES relied heavily on surface features.
With the advancement of deep learning since 2016, the technical paradigm shifted. Taghipour and Ng [2016] proposed an end-to-end AES using LSTM, demonstrating the effectiveness of neural approaches. Hussein et al. [2019] systematically classified AES research and identified three major challenges: human-likeness, deception resistance, and creativity assessment.
More recently, LLMs have been applied to AES. Naismith et al. [2023] evaluated discourse coherence using GPT-4, demonstrating the possibility of explainable AES. Xiao et al. [2024] showed that novice evaluators referring to LLM feedback achieved accuracy close to experts, redefining AES as human capability augmentation rather than complete automation. However, Bevilacqua et al. [2025] reported that LLM-based AES tends to rate AI-generated texts higher than human texts, revealing a fundamental bias problem. Zhou et al. [2026] demonstrated that LLMs can serve as effective tools for early prediction of student performance, further expanding the role of LLMs in educational assessment.
Throughout this evolution, AES research has consistently focused on predicting text quality rather than evaluating originality of argumentation. As Dıklı [2006] noted, creativity and critical thinking assessment remain recognized but unaddressed challenges. AROA addresses this gap by introducing argumentative originality as an evaluation axis independent of quality.
@@ 2.2 Argumentation Mining
Argumentation mining automatically identifies and extracts argument structures from natural language text. Lawrence and Reed [2020] provided a comprehensive survey covering claim identification, support and rebuttal relation extraction, and mapping to argumentation schemes. The Toulmin model [Toulmin, 2003], which structures argumentation through six elements (Claim, Grounds, Warrant, Backing, Qualifier, and Rebuttal), is widely adopted as the theoretical foundation for argumentative essay analysis.
Stab and Gurevych [2014] initiated argumentation mining research targeting student essays by proposing an annotation scheme for claims, premises, and support/attack relations. Wachsmuth et al. [2016] extended this direction from structure extraction to quality assessment by evaluating argument organization and strength. In educational applications, Nguyen and Litman [2018] integrated argument features into AES and demonstrated improved accuracy in persuasive essay scoring, while Wambsganss et al. [2022] developed a system that visualizes argument structures and provides feedback to promote learners’ self-correction.
Recent LLM-based approaches have further advanced the field. Li et al. [2024] proposed a multi-task benchmark revealing that LLMs show significant performance degradation when performing chained argument understanding. Shi et al. [2026] provided a comprehensive survey of computational argumentation, organizing the field into argumentation mining, assessment, and generation, yet originality assessment of argument content remains outside the scope of existing frameworks.
However, existing research evaluates argument quality (structural completeness, organization, persuasiveness) rather than originality of argument content. AROA utilizes argumentation mining techniques to extract argument elements and then quantifies their semantic rarity through density estimation, presenting a direction of evaluating originality rather than quality.
@@ 2.3 Originality and Creativity Assessment
Originality and creativity assessment has a long tradition in psychology and education. Runco and Yoruk [2014] defined creativity as the combination of novelty and usefulness, with divergent thinking tests, particularly the Alternative Uses Test (AUT), serving as the representative measurement paradigm.
Traditional originality assessment relied on human raters subjectively judging response rarity, but automatic evaluation based on semantic distance has since become mainstream. Beaty et al. [2022] proposed evaluating originality by calculating semantic distance using sentence embeddings, establishing the standard approach for automatic creativity assessment. Patterson et al. [2023] verified the multilingual validity of this approach across 12 languages and over 6,500 participants.
LLM-based creativity assessment has also advanced. DiStefano et al. [2025] fine-tuned LLMs for metaphor generation and achieved agreement with human evaluation exceeding semantic distance methods. Valueva et al. [2024] reviewed automatic scoring methods for divergent thinking tests, noting that while LLMs are approaching human-level evaluation, the opacity of judgment rationales remains a challenge. Runco et al. [2025] analyzed AI-generated divergent thinking responses, showing that AI generates numerous but often irrelevant and redundant ideas.
These studies have succeeded in evaluating novelty of short responses in divergent thinking tests. However, no research evaluates argumentative originality in long texts such as essays, specifically how rare argument elements (claims, evidence, counterarguments) are relative to a reference corpus. AROA extends the concept of rarity as originality, established in divergent thinking research, to argumentative essays by integrating rarity-based evaluation with argumentation mining across four complementary components.